OpenStudy (anonymous):

what is 3 times the square root of 10

myininaya (myininaya):

\[3 \sqrt{10}\]

myininaya (myininaya):

this is simplified

OpenStudy (turingtest):

He sounds crazy but he's right

myininaya (myininaya):

she* lol

OpenStudy (turingtest):

sorry again :(

myininaya (myininaya):

for what

OpenStudy (anonymous):

(-2)^4+ 3 square root (120-20)/4^3+ abs value 5(-1)

OpenStudy (turingtest):

I figure it comes of as sexism to assume anyone's a male is all.

OpenStudy (turingtest):

@kvargas you need to try to use parentheses better or utilize the "equation" button on the bottom left of your input box

OpenStudy (turingtest):

\[(-2)^4+3\sqrt{120-20}/4^3+\left| 5(-1) \right|\] is this correct?